Hi i need help solving this equation
109.30= 10/(1+i)+110/(1+i)^2
i= 0.05 but how do we get to the i value? Can you please show step by step method. Thank you.

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



In the first place, is NOT equal to . It is very close to that number but close is not equal and therefore it is dead wrong to use the equals sign.

Let then:



Multiply by



Multiply by 10 to eliminate the decimal and then use the quadratic formula to solve for



But so:



I'm going to go way out on a limb here and assume that you are looking for a non-negative value, so:



John

My calculator said it, I believe it, that settles it


Answer by stanbon(75887)

Hi i need help solving this equation
109.30= 10/(1+i)+110/(1+i)^2
i= 0.05 but how do we get to the i value? Can you please show step by step method.
-----
Multiply thru by (1+i)^2 to get::
110 + 10(1+i) = 109.3(1+i)^2
------
Let w = 1+i
-------
Solve for "w"::
109.3w^2 - 10w - 110 = 0
----
Use the Quadratic Formula to solve for "w"
w = -0.9585 or w = 1.05
----
Solve for "i"::
1+i = -0.95785 or 1+i = 1.05
Ans: i = -1.95785 or i = 0.05
----
Cheers,
Stan H.
----------